A child's bank contains $2.24 in pennies and nickels. If the number of pennies is 40 less than 3 times the number of nickels, then how many pennies are in the bank?
Answer by stanbon(75887)

A child's bank contains $2.24 in pennies and nickels. If the number of pennies is 40 less than 3 times the number of nickels, then how many pennies are in the bank?
----------------
Value Equation: p + 5n = 224 cents
Quantity Eq:::: p = 3n-40
----------------------------------------
Substitute to solve for "n":
3n-40 + 5n = 224
8n = 264
n = 33 (number of nickels)
---
Substitute into p = 3n-40 to solve for "p":
p = 3*33-40 = 99-40 = 59 (# of pennies)
=============================================
Cheers,
Stan H.